OpenStudy (anonymous):

please help on the one with the n-4_> 6

OpenStudy (anonymous):

I'm guessing that you mean \(n-4 \geq 6\). This is easy, just add 4 to both sides \[ n-4+4 \geq 6+4\]\[n \geq 10\]

OpenStudy (anonymous):

yes that is what i was looking for

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

If that was the equation and no extra information, then it is usual that it is just a simplification. So yes :)